OpenStudy (anonymous):

One cubic centimeter (1.0 cm3) of water has a mass of 1.0 10-3 kg. (a) Determine the mass of 1.0 m3 of water

OpenStudy (fifciol):

1.0 cm3 - 1.0 10-3 kg 1.0 m3 - x x* 1.0 cm3 = 1.0 10-3 kg m3 x* 1.0 (10-2m)3 = 1.0 10-3 kg m3 x=10-3/10-6 x=10^3 kg